OpenStudy (anonymous):

If f(x) = 3x + 2 and g(x) = x2 + 1, which expression is equivalent to (g0f) (x)

OpenStudy (anonymous):

to what ?

OpenStudy (anonymous):

If f(x) = 3x + 2 and g(x) = x2 + 1, which expression is equivalent to (GoF) (x)

OpenStudy (academicgurusinc):

Note: I am assuming that by g(x)= x2 + 1 you means g(x) = x^2 + 1 instead of (x*2) +1 gof(x) simply means that you will be replacing the "x" in in g(x) with f(x). Since g(x) = x^2 + 1, then gof(x) = (3x + 2)^2 + 1 = (3x + 2)(3x +2) +1 = 9x^2 + 6x + 6x +4 +1 = 9x^2 + 12x + 5 We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)